OpenStudy (anonymous):

let v1=(3,5) and v2=(-4,7) Compute the unit vectors in the direction of |v1| and |v2|????

OpenStudy (anonymous):

unit vector v1= \[\frac{ v1 }{ ||v1|| }\] just compute. that's it

OpenStudy (anonymous):

So you would put (3,5) on top and what on the bottom then?

OpenStudy (anonymous):

sorry, I messed up the denominator, let me redo

OpenStudy (anonymous):

Is it the v2 on the bottom then?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{ <3,5> }{ \sqrt{3^2 + 5^2} }= \frac{ <3,5> }{ \sqrt{34} }= <\frac{ 3 }{ \sqrt{34} },\frac{ 5 }{ \sqrt{34} }>\]

OpenStudy (anonymous):

no, we just use algebraic way to get unit vector

OpenStudy (anonymous):

Ok, So how would we compute them with v1 and v2 not with the unit vectors

OpenStudy (anonymous):